OpenStudy (anonymous):

Completely factor 2xy-4x+8y-16

OpenStudy (anonymous):

2xy - 4x + 8y - 16 = 2x(y - 2) + 8(y - 2) = (2x + 8)(y - 2) = 2(x + 4)(y - 2)

OpenStudy (anonymous):

hope this helps

OpenStudy (anonymous):

Alotttt